OpenStudy (anonymous):

Find the equations of the horizontal asymptotes and the vertical asymptotes of f(x): f(x)= (7x^2+7x+10)/(2x^+11x-63) thanks

OpenStudy (anonymous):

on the denominator do you mean 2x^2

OpenStudy (anonymous):

yes, sorry about that.

OpenStudy (anonymous):

okay the vertical asymptotes are the lines x=a where a is a zero of the denominator. In this case the denominator factors to (x+9)(2x-7), therefore the vert. asymptotes occur at -9 and 2/7

OpenStudy (anonymous):

horizontal asy is y=7/2

OpenStudy (anonymous):

Now, as for the horizontal, it can be determined by the leading terms of both the numerator and denominator. More, specifically, i this case the exponents are the same (7x^2 and 2x^2), therefore we take only the ratio of the coefficients, thus the hor. asymptote is a y=7/2, as ictrees so eloquently put it

OpenStudy (anonymous):

HA = 7/2 VA = -9 and 7/2. thanks guys.